Pipe A can fill a tank in 14 hours. Pipe B can empty a tank in 10 hours. Miscommunication between Emilio and Stephen causes Emilio to release water into Pipe A at the same time as Stephen opens Pipe B to empty the tank. With both pipes open to the tank, how long does it take to empty the tank?
Answer by stanbon(75887)

Pipe A can fill a tank in 14 hours. Pipe B can empty a tank in 10 hours. Miscommunication between Emilio and Stephen causes Emilio to release water into Pipe A at the same time as Stephen opens Pipe B to empty the tank. With both pipes open to the tank, how long does it take to empty the tank?
----
Pipe A rate = 1/14 job/hr
Pipe B rate = 1/10 job/hr
------
Together rate = 1/x job/hr
=================================
Note: The job is to empty the tank, so Pipe A has a negative rate.
Equation:
rate - rate = together rate
-1/14 + 1/10 = 1/x
-10x + 14x = 14*10x
4x = 140x
x= 35 hrs (time to emptythe tank working together)
============
Cheers,
Stan H.